what are original dimensions- The length of a rectangle is three times its width. If the width were increased by 4 and the length remained the same, the resulting rectangle would have an area of 231 square inches. Find the demensions of the ORIGINAL rectangle.

Answer by mananth(16949)

what are original dimensions- The length of a rectangle is three times its width. If the width were increased by 4 and the length remained the same, the resulting rectangle would have an area of 231 square inches. Find the demensions of the ORIGINAL rectangle.

let the width be x
length will be 3x
width increased by 4 = x+4

Length = 3x
3x(x+4)= 231
3x^2 + 12x =231
3x^2 +12x-231=0
3(x^2+4x-77) =0
3(x+11)(x-7)=0
x=-11 OR x =7
the width cannot be negative
so width = 7
length = 3x = 3*7 =21
7 by 21 are the dimensions of the original rectangle